OpenStudy (anonymous):

In a game with eight contestants, half the contestants are eliminated each round. The winners of the first round move to the second round. The winners of the second round move to the third round. The winner of the third round wins the whole game. First place and second place medals are awarded to the two contestants who made it into the third round. How many different arrangements of contestants can be awarded first place and second place medals? 16?

OpenStudy (turingtest):

i think it's 8x7=56 but I'm not sure...

OpenStudy (anonymous):

they do have a 56 option but how'd you get that?

OpenStudy (anonymous):

8!/2!(8! - 2!)?

OpenStudy (turingtest):

If the first medal can go to one of the eight, that only leaves seven contestants left who could receive the 2nd medal. in other words the odd of winning the first place=1/8, the odds of now getting the second are now 1/7 since one person is out of the competition., (1/8)(1/7)=1/56 where 56 are all possible configurations. 8!/6! i guess

OpenStudy (anonymous):

ok I get it kinda that was last one thanks